Braden has $20 to spend on video game rentals (R), priced at $4 per game, and candy bars (C), priced at $1 each.

The equation of Brandon’s budget constraint (if R is on the X axis) is:

a.

C = 20 – 0.25R

b.

C = 5 – 0.25R

c.

C = 20 – 4R

d.

C = 5 – 4R
Answer #1

Budget constraint : Pr*R + Pc*C = M

4R + C = 20

C = 20 - 4R

Option c is correct